. At a pressure of 1.5 atm and 42.2 °C, a certain gas has a volume of 340.0 mL. What will be the volume of this gas under STP.

 


1
Expert's answer
2021-04-07T11:31:53-0400

Solution:

STP conditions (a temperature of 273.15 K and a pressure of 1 atm).

Hence,

P1 = 1.5 atm

V1 = 340.0 mL

T1 = 42.2°C + 273.15 = 315.35 K

P2 = 1.0 atm

V2 = unknown

T2 = 273.15 K


The Combined Gas Law can be used:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (1.5 atm × 340.0 mL × 273.15 K) / (1.0 atm × 315.35 K) = 441.75 mL = 441.8 mL

V2 = 441.8 mL


Answer: The volume of this gas will be 441.8 mL.